# Coulomb Interaction-Stabilized Isolated Narrow Bands with Chern Numbers in Twisted Rhombohedral Trilayer-Bilayer Graphene

### v1

@@ Abstract
Fractionalized phases that emerge from doping a quantum anomalous Chern insulator are highly-sought after because they may host quasiparticles with non-Abelian exchange statistics, importantly in lattice systems without a magnetic field [1, 2, 3, 4, 5, 6, 7]. These exotic quasiparticles are fascinating on their own right [8], but much of the excitement surrounding them concerns their potential application in topological quantum computation [9]. Within the past two years, these fractional Chern insulating (FCI) states have finally been experimentally measured in two classes of materials: twisted MoTe2 bilayers without displacement field [10, 11, 12, 13, 14, 15, 16, 17, 18] and rhombohedral graphene multilayers aligned with hexagonal boron nitride (hBN) with displacement field [19, 20, 21]. In twisted MoTe theoretical calculations suggest that the fractional states emerge from partial filling of a valence band with that exists already in non-interacting band structure modeling [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In rhombohedral graphene-hBN heterostructures, fractional states have only been observed at displacement fields V/nm where the conduction electrons are pushed away from the substrate [19, 20, 21]. In this regime, the non-interacting conduction bands are essentially those of pristine rhombohedral graphene multilayer with minimal influence from the hBN substrate [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. That is, the conduction band where electrons would populate is nearly dispersionless but it has no discernible gap to the next higher-energy band [44, 45, 46]. However, upon including electron-electron interactions, at least at the mean-field level, the narrow band decouples from the other energy bands, generating a mean-field gap meV, and carrying a Chern number [32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. This band remains more or less flat at the mean-field level, and is thus, conductive to further strong interaction effects at partial filling. These graphene-based platforms offer the insight that non-interacting band structures can only predict a subset of possible topological materials that can fractionalize, i.e. they can predict twisted MoTe2 but not graphene-hBN moiré systems.
@@
Spectrally, the band structures for both the antiparallel and parallel configurations are qualitatively similar. However, their differences are encoded in the quantum geometry of the Bloch states. For cases where the conduction band is fully gapped666Fully gapped here means that it is gapped both above and below., we calculate its Chern number and associated quantum metric violation defined as where is the quantum metric and is the Berry curvature [100, 101, 102, 103, 104]. The quantity is a proxy for the propensity of a band to become an FCI once doped. It is based on analogy to the quantum geometry of the zeroth Landau level [78, 79, 80]. We shall use as a possible FCI diagnostic. The smaller it is, the better for FCI’s. For comparison, the trace condition violations calculated for systems that have been experimentally shown to host FCI states are comparable to our values. For graphene pentalayer on hBN, [34, 32, 37, 33]; for twisted MoTe [105, 22, 29]. The Chern characterization of the antiparallel and parallel configurations in space is shown in Fig. 1(b,d). We find that the antiparallel configuration hosts a large region where In this region, the smallest trace condition violation is This coincides with the where the bandwidths are comparatively small ( meV) and band gaps relatively large meV777There are also regions where Since we are interested in higher Chern numbers in this work, we do not analyze these regions here. However, it is perhaps worth mentioning that for these regions. The bandwidth of the region is relatively small ( meV), but the bandwidth of the region is quite large ( meV).. The parallel configuration is even more exotic in that it hosts both a and a region. The region here features a slightly smaller trace condition violation, with at best, compared to the antiparallel configuration. On the other hand, the region has quite small However, the bandwidths of the phase are generally larger ( meV) than the phase ( meV). All in all, we find that for both the parallel and antiparallel configurations, there are regions with nonzero Chern numbers where relatively small bandwidths, large band gaps, and small coexist. We mark these regions with dashed lines in Fig. 1(b,d). From each region, we display a representative band structure in Fig. 1(c,e) showing the narrow band of interest polarized primarily to layer 1. We also show three band structures in Fig. 1(f,g,h) that seemingly do not look especially conductive to interaction effects, but yet they will turn out to be important once interactions are included.
@@
Next, we analyze the stability of these phases as the Coulomb interaction is modified. This is an important consideration because the strength of electron-electron interactions can be experimentally controlled by dielectric engineering (changing ) and gate distance (changing ). For this purpose, we park ourselves at a particular and sweep For demonstration, we choose the optimal combinations of as diagnosed by in the preceding analysis. The results are shown in Fig. 4. In general, all of these states seem quite insensitive to changing the gate distance The dependence on is more intricate. In Fig. 4(a), we assess the stability of the antiparallel configuration for a state. As is clear, this state is robust for At the highest (i.e. the weakest Coulomb strength), this state is out-competed by the symmetric metallic state. The value of appears optimal for this state and lies in the range of that are typically used to . remains relatively unchanged as and are varied. For a state in the parallel configuration, we again observe that there is a finite range of where this phase exists. For weaker Coulomb interactions (larger ), this state is outcompeted by the symmetric metallic phase too. For smaller , this state merges into a state. The state in the parallel configuration is much more robust than the other two states. It persists essentially in the entire domain of for our scan. The optimal dielectric constant occurs around We also show extensive phase diagrams for in Ref. [98].

### v2

@@ Abstract
Fractionalized phases that emerge from doping a quantum anomalous Chern insulator are highly-sought after because they may host quasiparticles with non-Abelian exchange statistics, importantly in lattice systems without a magnetic field [1, 2, 3, 4, 5, 6, 7]. These exotic quasiparticles are fascinating on their own right [8], but much of the excitement surrounding them concerns their potential application in topological quantum computation [9]. Within the past two years, these fractional Chern insulating (FCI) states have finally been experimentally measured in two classes of materials: twisted MoTe2 bilayers without displacement field [10, 11, 12, 13, 14, 15, 16, 17, 18] and rhombohedral graphene multilayers aligned with hexagonal boron nitride (hBN) with displacement field [19, 20, 21]. In twisted MoTe theoretical calculations suggest that the fractional states emerge from partial filling of a valence band with that exists already in non-interacting band structure modeling [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In rhombohedral graphene-hBN heterostructures, fractional states have only been observed at displacement fields V/nm where the conduction electrons are pushed away from the substrate [19, 20, 21]. In this regime, the non-interacting conduction bands are essentially those of pristine rhombohedral graphene multilayer with minimal influence from the hBN substrate [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. That is, the conduction band where electrons would populate is nearly dispersionless but it has no discernible gap to the next higher-energy band [44, 45, 46]. However, upon including electron-electron interactions, at least at the mean-field level, the narrow band decouples from the other energy bands, generating a mean-field gap meV, and carrying a Chern number [32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. This band remains more or less flat at the mean-field level, and is thus, conducive to further strong interaction effects at partial filling. These graphene-based platforms offer the insight that non-interacting band structures can only predict a subset of possible topological materials that can fractionalize, i.e. they can predict twisted MoTe2 but not graphene-hBN moiré systems.
@@
Spectrally, the band structures for both the antiparallel and parallel configurations are qualitatively similar. However, their differences are encoded in the quantum geometry of the Bloch states. For cases where the conduction band is fully gapped666Fully gapped here means that it is gapped both above and below., we calculate its Chern number and associated quantum metric violation defined as where is the quantum metric and is the Berry curvature [100, 101, 102, 103, 104]. The quantity is a proxy for the propensity of a band to become an FCI once doped. It is based on analogy to the quantum geometry of the zeroth Landau level [78, 79, 80]. We shall use as a possible FCI diagnostic. The smaller it is, the better for FCI’s. For comparison, the trace condition violations calculated for systems that have been experimentally shown to host FCI states are comparable to our values. For graphene pentalayer on hBN, [34, 32, 37, 33]; for twisted MoTe [105, 22, 29]. The Chern characterization of the antiparallel and parallel configurations in space is shown in Fig. 1(b,d). We find that the antiparallel configuration hosts a large region where In this region, the smallest trace condition violation is This coincides with the where the bandwidths are comparatively small ( meV) and band gaps relatively large meV777There are also regions where Since we are interested in higher Chern numbers in this work, we do not analyze these regions here. However, it is perhaps worth mentioning that for these regions. The bandwidth of the region is relatively small ( meV), but the bandwidth of the region is quite large ( meV).. The parallel configuration is even more exotic in that it hosts both a and a region. The region here features a slightly smaller trace condition violation, with at best, compared to the antiparallel configuration. On the other hand, the region has quite small However, the bandwidths of the phase are generally larger ( meV) than the phase ( meV). All in all, we find that for both the parallel and antiparallel configurations, there are regions with nonzero Chern numbers where relatively small bandwidths, large band gaps, and small coexist. We mark these regions with dashed lines in Fig. 1(b,d). From each region, we display a representative band structure in Fig. 1(c,e) showing the narrow band of interest polarized primarily to layer 1. We also show three band structures in Fig. 1(f,g,h) that seemingly do not look especially conducive to interaction effects, but yet they will turn out to be important once interactions are included.
@@
Next, we analyze the stability of these phases as the Coulomb interaction is modified. This is an important consideration because the strength of electron-electron interactions can be experimentally controlled by dielectric engineering (changing ) and gate distance (changing ). For this purpose, we park ourselves at a particular and sweep For demonstration, we choose the optimal combinations of as diagnosed by in the preceding analysis. The results are shown in Fig. 4. In general, all of these states seem quite insensitive to changing the gate distance The dependence on is more intricate. In Fig. 4(a), we assess the stability of the antiparallel configuration for a state. As is clear, this state is robust for At the highest (i.e. the weakest Coulomb strength), this state is out-competed by the symmetric metallic state. The value of appears optimal for this state and lies in the range of that are typically used [112]. remains relatively unchanged as and are varied. For a state in the parallel configuration, we again observe that there is a finite range of where this phase exists. For weaker Coulomb interactions (larger ), this state is outcompeted by the symmetric metallic phase too. For smaller , this state merges into a state. The state in the parallel configuration is much more robust than the other two states. It persists essentially in the entire domain of for our scan. The optimal dielectric constant occurs around We also show extensive phase diagrams for in Ref. [98].
@@
Note added: After this work had been posted to the arXiv, Ref. [115] reported experimental observations of integer Chern insulating states in at for displacement fields that drive conduction electrons towards the outer layer of the trilayer substack. These observations are entirely consistent with our predictions of stable bands in the parallel configuration for negative and angles